Documenta Mathematica

Band 14, 2009

Jaya NN Iyer, Stefan Müller–Stach Chow–Künneth Decomposition

for Some Moduli Spaces  1–18

Dennis Gaitsgory and David Nadler Hecke Operators on Quasimaps

into Horospherical Varieties  19–46

Mark L. MacDonald Projective Homogeneous Varieties

Birational to Quadrics  47–66

iv